Find the distance between an electron and a proton if the electric force between them is 82.4 nN.
Answer #1

Both have same charge=q=1.6*10^-19C

kq^2/r^2=82.4*10^-9

9*10^9*q^2/r^2=82.4*10^-9

substituting values

r=.5287*10^-10m